A bag contains 4 RED beads, 5 BLUE beads, and 11 GREEN beads.
If a single bead is picked at random, what is the probability that the bead is RED or GREEN?
Answer #1

Given that a bag contains 4 RED beads, 5 BLUE beads, and 11 GREEN beads.

so, P(Red) = 4/(4+5+11) = 0.2

P(Blue) = 5/(4+5+11) = 0.25

and P(Green) = 11/(4+5+11) = 0.55

Since all are independent and mutually exclusive hence

and as we know P(A or B) = P(A) +P(B) - P(A)*P(B)

hence P( Red or Green) = P(Red) + P(Green) - 0

P( Red or Green) = 0.2 + 0.55

P( Red or Green) = 0.75